10. Budget balances and the national debt The following table lists federal expenditures, revenues, and GDP for a hypotheticaExpenditures A Revenues EXPENDITURES AND REVENUES (Percent of GDP) 1996 1998 2000 2002 YEAR 2004 2006In which years was the governments budget in deficit? Check all that apply. 0 О 1998 О2000 о 2002 o2004 0 О 2006 In 2000, th
Answer #1

(1)

Revenue as % of GDP = (Revenue / GDP) x 100%

Expenditure as % of GDP = (Expenditure / GDP) x 100%

Data table:

Year Revenue Expenditure GDP Revenue/GDP Expenditure/GDP
1998 1,722 1,653 8,747 20% 19%
2000 2,025 1,789 9,817 21% 18%
2002 1,853 2,011 10,470 18% 19%
2004 1,880 2,293 11,686 16% 20%
2006 2,407 2,655 13,178 18% 20%

Graph:

Expenditures Revenues EXPENDITURES AND REVENUES (Percent of GDP) 1996 1998 2000 2004 2006 2002 YEAR

(2) Budget was in deficit when Revenue < Expenditure, in years 2002, 2004 and 2006.

(3) In 2000, national debt decreased by $236 Billion by 242.03%.

National debt in year N = National debt in year (N - 1) + (Expenditure - Revenue)year N

In 1998, National debt = 1,653 - 1,722 = - 69

In 2000, National debt = - 69 + (1,789 - 2,025) = - 305

Decrease in national debt = - 69 - (- 305) = - 69 + 305 = 236

% Decrease in national debt = (236 / 69) - 1 = 3.4203 - 1 = 2.4203 = 242.03%.